Question 1.Susan can brew 5 gallons of root beer in an hour or she can make 4 pizzas in an hour. Becky can brew 7 gallons of root beer in an hour or she can make 5 pizzas in an hour. Susan's opportunity cost of making a pizza is gallons of root beer. Becky's opportunity cost of making a pizza is gallons of root beer. Enter numbers rounded to two decimal places.

Question 2 Bill and Fred bake cakes and pies. Bill's opportunity cost of baking 1 pie is 3 cakes. Fred's opportunity cost of baking 1 pie is 5 cakes. Who should bake the pie? Bill or Fred

Question 3 For both parties to gain from trade, the price at which they trade must lie between the two opportunity costs. True or False

Question 4 If the resources used to produce two goods are specialized, then the production possibilities frontier has a bowed­ out shape. True or False

Question 5 Two countries produce two goods and they each have 4 million hours of labor. Consider the table that shows the quantity produced per labor hour: If each country specializes in the good for which it has a comparative advantage then in total the two countries can together produce _______million shirts and _____million bushels of corn. Enter whole numbers.

Shirts Corn(Bushels)
Alpha 4 12
Beta 2 10

Question 8 A country should specialize in producing the goods for which it has the highest opportunity cost. True or False

Question 9 4 pts If a country has a comparative advantage in producing fish, then they must also have an absolute advantage in producing fish. True or False

Question 10 If there are constant opportunity costs of production then the production possibilities frontier is a straight line. True or False

Question 11 If John can produce a good in less time than Fred then he has an absolute advantage in producing that good. True or False

Question 12 Latasha can read 30 pages of economics in an hour. She can also read 20 pages of sociology in an hour. She spends 4 hours per day studying. Her opportunity cost of reading 30 pages of sociology is ______ pages of economics. Her opportunity cost of reading 15 pages of economics is ______ pages of sociology. Enter whole numbers.

Question 13 Suppose that in a week an American worker can produce 60 shirts or 3 computers and a Chinese worker can produce 60 shirts or 1 computers. The country that specializes in computers will trade 1 computer in exchange for at least shirts. The country that buys the computer will pay not more than shirts. Enter whole numbers.

Question 14 Consider the following production possibilities table that shows different combinations of two goods that can be produced with given resources:

Based on the table, which is true? Opportunity cost is constant OR Opportunity cost is increasing

Shirts Corn
A 0 20
B 10 15
C 20 10
D 30 5
E 40 0

Question 15 A country can produce either 25 pounds of fish or 7 bushels of corn. What is the opportunity cost of producing one bushel of corn? Enter a number rounded to two decimal places.

Question 16 England and Scotland both produce scones and sweaters. Suppose that an English worker can produce 30 scones per hour or 2 sweater per hour. Suppose that a Scottish worker can produce 40 scones per hour or 4 sweaters per hour. Who will produce sweaters? Scotland England

Question 17 Susan can brew 5 gallons of root beer in an hour or she can make 4 pizzas in an hour. Becky can brew 7 gallons of root beer in an hour or she can make 5 pizzas in an hour. Who has an absolute advantage in making pizza? Susan or Becky

Question 18 Bill and Fred bake cakes and pies. Bill's opportunity cost of baking 1 pie is 3 cakes. Fred's opportunity cost of baking 1 pie is 5 cakes. If both parties are to benefit from trade then we can expect 1 pie to sell for:

A) 4 cakes

B) 2 cakes

C) 6 cakes

Question 19 A country decides to specialize in wheat production. They now export wheat and import shirts. Which groups of people in this country benefit from trade? PICK all that are correct.

A) wheat consumers

B) wheat producers

C) shirt consumers

D) shirt producers

Question 20 When countries specialize in producing certain goods and then freely exchange those goods for other goods with different countries, what is the advantage?

A) Each country can consume at a point outside their production possibilities frontier.

B) All people in each country will be better off than they would be if they did not trade.

C) Each country can produce and consume at a point outside their production possibilities frontier.

D) Each country can produce at a point outside their production possibilities frontier.

Question 21 Susan can brew 5 gallons of root beer in an hour or she can make 4 pizzas in an hour. Becky can brew 7 gallons of root beer in an hour or she can make 5 pizzas in an hour. Who has a comparative advantage in making root beer? Becky or Susan

Question 22 Two countries can achieve gains from trade even if one of the countries has an absolute advantage in the production of all goods. True or False

Question 24 Consider the following production possibilities table that shows different combinations of two goods that can be produced with given resources:

Shirts Corn
A 0 20
B 10 15
C 20 10
D 30 5
E 40 0

When the country moves from point A to point B, the opportunity cost of producing 1 shirt is ______ bushels of corn.

When the country moves from point C to point D, the opportunity cost of producing 1 shirt is ______ bushels of corn. Enter numbers rounded to two decimals as necessary.

Question 23 If trade is beneficial for a country, it must be good for everyone in the country. True or False

Question 25 A country that has an absolute advantage in producing corn does not necessarily have a comparative advantage in producing corn. True or False
Answer #1

1.

For Susan, opportunity cost of making pizza = 5/4 = 1.25 gallons root beer

For Becky, opportunity cost of making pizza = 7/5 = 1.4 gallons of root beer

2.

Bill should bake pie and Fred should bake cakes. It is based on the relatively lower opportunity cost and subsequent comparative advantage.

3.

True

It will make better-off to both the parties after the trade.

4.

True

5.

For alpha, opportunity cost to produce 1 shirt = 12/4 = 3 bushels of corn

For Beta, opportunity cost to produce 1 shirt = 10/2 = 5 bushels of corn

So, Alpha will produce shirts due to lower opportunity cost and Beta will produce corn.

Total Shirts to be produced = 4 Million/4 = 1 Million shirts

Total corn to be produced = 4 Million/10 = .4 Million bushels of corn

Pl. repost other unanswered questions for their proper answers!